3. The profit function for a company is given by the following nonlinear equation:
P(x) = 4x2 – 200x − 25000 where P is the profit in dollars and x is the number of units sold.
3.1.1. Find the profit when 600 units are sold. (1 mark)
3.1.2. Determine the break-even level for this company. (3 marks)
1
Expert's answer
2019-11-20T10:46:25-0500

3.1.1."P(X)=4X^2 -200X-25000"

           P(600)=4*(600*600)-200(600)-25000

          P(600)=1440000 -120000 -25000

           P(600) =1295000

           Where p(600) is the profit when 600 units are sold.

3.1.2.  at P(x)=o

            4x^2 - 200x - 25000 = 0. Hence solving the quadratic equation;

           First divide the equation all through by 4;

          "X^2-50X-6250=0"

"X=25(1+\u221a11), X=25(1-\u221a11)"

           X=107.92 OR X=-57.92